how many one-fifths are there in 4 and 2/3 ?
one-fourth a number, raised to the power, gives 27/64. what is the number ?
1/x-1/y=1/z.if x=2 and y=3, what is z ?

Answer by JBarnum(2146)

4 times 3 plus 2 = 14
14/3 times 5/5= 70/15
1/5=3/15
70/3=23.3333
so 23 one fifths are in 4 and 2/3
________________
raised to what power??
___________
fill in the variables
1/2-1/3=3/6-2/6=1/6 so z has to be 6